# The Internal Model Principle of Time-Varying Optimization

### v2

@@ Abstract
Time-varying optimization problems are central to many engineering applications, where performance metrics and system constraints evolve dynamically with time. Several algorithms have been proposed to address these problems; a common characteristic among them is their implicit reliance on knowledge of the optimizers’ temporal variability. In this paper, we provide a fundamental characterization of this property: we show that an algorithm can track time-varying optimizers if and only if it incorporates a model of the temporal variability of the optimization problem. We refer to this concept as the internal model principle of time-varying optimization. Our analysis relies on showing that time-varying optimization problems can be recast as output regulation problems and, by using tools from center manifold theory, we establish necessary and sufficient conditions for exact asymptotic tracking. As a result, these findings enable the design of new algorithms for time-varying optimization. We demonstrate the effectiveness of the approach through numerical experiments on both synthetic problems and the dynamic traffic assignment problem from traffic control.
@@ 1 Introduction
Time-varying optimization problems play a central role in several scientific domains, as they underpin many important contemporary engineering problems. Examples include training in Machine Learning [1, 2], dynamic signal estimation in Signal Processing [3], trajectory tracking in Robotics [4], system optimization in Industrial Control [5], and much more. Historically, discrete-time algorithms for time-varying optimization have been proposed and studied first since they emerged as a natural extension of their time-invariant counterparts, allowing for cost functions and constraints that may change over time [6]. These approaches build on the classical perspective on mathematical optimization, which seeks to construct methods to determine optimizers and consist of iterative procedures implemented on digital devices. In recent years, however, numerous optimization algorithms have been formulated and studied as continuous-time processes [7, 8, 9], mainly due to the wide availability of tools for Ordinary Differential Equations (ODEs) that can facilitate their analysis [9].
Motivated by these recent developments, in this paper, we study time-varying convex optimization problems and focus on the use of continuous-time dynamics to track exactly optimal solutions. Several approaches have already been developed for this purpose [10, 11, 12, 13], yet all these techniques implicitly require full knowledge of the temporal variability of the problem [13]. Unfortunately, in most practical applications, having such knowledge is impractical, either because the temporal variability enters the optimization in the form of exogenous disturbances that are unknown and cannot be measured (see, e.g., [14, 15]), or simply because it is unrealistic to ask for a noiseless model of how the problem changes with time. Departing from this, in this paper we pose the following question: is it possible to track (exactly and asymptotically) a minimizer of a time-varying optimization problem without any knowledge of the temporal variability of the optimization? Interestingly, we prove a fundamental result showing that tracking can be achieved if and only if the temporal variability of the problem can be ‘observed’ by the algorithm, and the latter incorporates a suitably reduplicated model of such a variability. We refer to this conclusion as the internal model principle of time-varying optimization, akin to its control-theoretic counterpart [16, 17]. Our approach relies on reinterpreting the optimization algorithm design as a nonlinear, multivariable regulation problem [18], and our analysis uses tools from center manifold theory [19, 17]. Fig. 1 illustrates the architecture of the gradient-feedback design scheme studied in this paper.
@@
Figure 1: Architecture of the gradient-feedback design scheme studied in this work. An optimization algorithm is to be designed (bottom block), having access only to gradient evaluations of the loss function to be minimized (top right block), and generating a sequence of exploration points $x(t)$ at which the gradient shall be evaluated. The loss function to be optimized varies with time, where the temporal variability $\theta(t)$ is assumed to be unmeasurable and generated by an exosystem (top left block). Shaded blocks emphasize the presence of dynamics.
@@
Related works. The literature on methods for time-varying optimization is mainly divided into two classes of solutions. The first class consists of methods that do not utilize any model of the temporal variability of the problem [20, 21, 22, 23]; instead, they seek to solve a sequence of static problems. Several established approaches belong to this class, including the online gradient descent method [24] and the online Newton step algorithm [25] – see [26] and references therein. Because the temporal variability of the problem is unknown (or ignored), algorithms in this class can make a decision only after each variation has been observed, thus incurring a certain ‘regret.’ Mathematically, these techniques are capable of reaching only a neighborhood of an optimizer [13], and exact tracking is out of reach in general for these approaches. Moreover, suitable assumptions need to be made to provide convergence guarantees for these methods; we refer to [13, 26] and references therein for a detailed discussion. In contrast, the second class of methods uses a model of the temporal evolution of the problem to exactly track the optimal trajectories [27, 2, 28, 29]. Particularly celebrated is the prediction-correction algorithm (see [10], the recent work [11], and the discrete implementations [30]), whereby at each time a prediction step is used to anticipate how the optimizer will evolve over time, and a correction step is used to seek a solution to each instantaneous optimization problem. We refer to [13] for a recent overview of the topic. Recent years have witnessed a growing interest in this problem: [4] uses contraction to study these methods; [31] uses sampling to estimate the temporal variability of the problem; a recent survey has appeared in [32]; constrained optimization problems are studied in [33]; recent works [34, 35, 36, 37] have modified these methods to achieve fixed-time convergence. With respect to this body of literature, which assumes precise knowledge of the temporal variability of the problem (see detailed discussion in Section 2.1), our focus here is to answer the following fundamental question: what is the least-restrictive information that is needed to track, asymptotically with zero error, the optimizer of a time-varying problem? To the best of the authors’ knowledge, a rigorous answer to this question is missing in the literature.
@@
In this work, we recast the problem of designing optimization algorithms for time-varying optimization as a nonlinear output regulation problem. This connects our work with the literature on output regulation, which is a well-established area of research. Initial works can be traced back to the 1970s [16, 38, 18] focusing on linear systems, and later extended to nonlinear systems using both local [39, 40] and global [41] approaches for the analysis. In recent years, the field has received new attention using modern methods, see, e.g., [42, 43, 44] and the recent tutorial [45]. To the best of our knowledge, this is the first work in the literature that establishes a connection between output regulation and optimization algorithms design.
@@
Organization. Section 2 presents the problem of interest, Section 3 the parameter-feedback problem (where the temporal variability of the problem can be explicitly measured), and Section 4 the gradient-feedback problem (where the algorithm has access only to first-order functional evaluations). Section 5 discusses the tracking accuracy in relationship to the fidelity of the internal model, Section 6 presents extensions to constrained optimization and discrete-time algorithms, Section 7 validates numerically the results, and Section 8 illustrates our conclusions. Finally, in the Appendix, we summarize basic facts on center manifold theory used in the paper.
[…]
@@ 2.1 Optimization objectives
$\min_{x\in{}^{n}}f(x,\theta(t)),$ (1)
where $t\in\mathbb{R}_{\geq 0}$ denotes time and $f:{}^{n}\times\Theta\rightarrow\real,$ $\Theta\subseteq{}^{p},$ is a loss function that is parametrized by the time-varying parameter vector $\theta:\mathbb{R}_{\geq 0}\rightarrow\Theta.$ We make the following standard assumptions on the loss.
@@
The map $x\mapsto f(x,\theta)$ is convex and $x\mapsto\nabla_{x}f(x,\theta)$ is Lipschitz continuous in n, for each $\theta\in\Theta.$square
@@
There exists a smooth (i.e., $C^{\infty}$) vector field $s:\Theta\rightarrow{}^{p}$ and $\theta(0)\in\Theta$ such that the parameter vector $\theta(t)$ satisfies
@@
For simplicity of the presentation, we assume that $\Theta$ is some neighborhood of the origin of p. We put no restrictions on the size of this neighborhood (which is, e.g., allowed to be the entire space $\Theta={}^{p}$), and thus on the size of $\theta(t)$ nor on its temporal variation. Moreover, there is no restriction with asking that $\Theta$ contains the origin because, if $\theta(t)$ takes values in the neighborhood of any other point, such a point can be shifted to the origin through a time-invariant change of variables without altering the solutions of (1).
@@
In what follows, we say that $x^{\star}(t),$ with $x^{\star}:\mathbb{R}_{\geq 0}\rightarrow{}^{n}$, is a critical trajectory of (1) if it satisfies:
@@
In our analysis, we will use a critical point with $\theta$ at rest, which is a constant vector $x^{\star}_{\circ}\in{}^{n}$, defined implicitly as:
@@
where $\eta>0$ models the algorithm’s step size. It is known [48] that this class of algorithms is capable of computing (asymptotically) critical points of (1) only when $\theta(t)$ is a constant signal; in all other cases, the algorithm converges to a neighborhood of a critical point and convergence is inexact. We refer the reader to Remark 11 for an insightful interpretation of this fundamental limitation. $\square$
@@
where $\nabla_{xx}f(x,\theta)$ is the Hessian of $f(x,\theta)$, and $\nabla_{x\theta}f(x,\theta)$ is the partial derivative of $\nabla_{x}f(x,\theta)$ with respect to $\theta$. Under strong convexity assumptions, this algorithm is known to converge exponentially, with zero asymptotic error, to a critical trajectory [10, 11]. Variations of this algorithm have also been proposed to achieve fixed-time or finite-time convergence [34]. We conclude by noting that implementing (8) requires (i) the cost to be twice continuously differentiable and strongly convex, (ii) knowledge of the maps $s(\theta)$, $\nabla_{x}f(x,\theta)$, $\nabla_{xx}f(x,\theta)$, and $\nabla_{x\theta}f(x,\theta)$, and (iii) knowledge of $\theta(t).$square
@@ 2.2 Problem statement
In line with the literature [26, 10, 11, 34, 12], we focus on gradient-type algorithms for solving (1), which are algorithms that have access to first-order oracles of the cost; that is, functional evaluations of the map:
@@
In the optimization literature [26], functional evaluations of (9) are generally obtained through two main approaches: (i) the algorithm has access to both the analytical form of $\nabla_{x}f(x,\theta)$ and $\theta(t)$ at each time $t$ (either because this signal is known or measurable), or (ii) an oracle returns evaluations of (9) (computed, e.g., through numerical differentiation). These alternatives will be discussed more in detail in (O1)–(O4), shortly below. $\square$
We let the optimization algorithm be described by a dynamic internal state $z(t),$ taking values on an open subset $\mathcal{Z}\subseteq{}^{n_{c}},n_{c}\in\mathbb{N}_{>0}$. The optimization algorithm generates a sequence of points $x(t)\in{}^{n}$ (called exploration signal), and has access to functional evaluations of (9) at these points (called gradient feedback signal):
@@
where $F_{c}:\mathcal{Z}\times{}^{n}\rightarrow{}^{n_{c}}$ and $G_{c}:\mathcal{Z}\rightarrow{}^{n}$ are functions to be designed. In the remainder, we refer to (10) as a dynamic gradient-feedback optimization algorithm. The architecture of the presented gradient-feedback algorithm is illustrated in Fig. 1. For simplicity of the presentation, we will require that $F_{c}(z,y)$ and $G_{c}(z)$ are such that [^3] Notice that this is without loss of generality since $F_{c}(z,y)$ and $G_{c}(z)$ are to be designed and $x_{\circ}^{\star}$ is known through (6).:
00 0 $=F_{c}(z_{\circ}^{\star},0)\qquad\text{and}\qquad x_{\circ}^{\star}=G_{c}(z_{\circ}^{\star}),$ (11)
for some locally unique $z_{\circ}^{\star}\in{}^{n_{c}}$. Together with (6), this ensures that $z_{\circ}^{\star}$ is an equilibrium of (10) when the gradient feedback signal $y(t)$ is equal to zero.
@@
Table 1: Summary of required knowledge and known guarantees by various established approaches. The symbol ()($\checkmark$)( ✓ ) is used when (O1) follows from (E1) and (O2). See Remarks 4, 5, and 7 for discussions.
@@ 3 The parameter-feedback problem
where $H_{c}:\Theta\rightarrow{}^{n}$ is a mapping to be designed. Because of the explicit dependence on $\theta(t),$ we will refer to (14) to as a parameter-feedback optimization algorithm. In analogy with (11), for simplicity of the presentation, in this section, we will impose that $H_{c}(\theta)$ satisfies:
@@
For the algorithm (14) considered in this section, Problems 2.2 and 1 are reformulated as follows.

### v3

@@ Abstract
Time-varying optimization problems are central to many engineering applications, where performance metrics and system constraints evolve dynamically with time. Several algorithms have been proposed to address these problems; a common characteristic among them is their implicit reliance on knowledge of the optimizers’ temporal variability. In this paper, we provide a fundamental characterization of this property: we show that an algorithm can track time-varying optimizers if and only if it incorporates a model of the temporal variability of the optimization problem. We refer to this concept as the internal model principle of time-varying optimization. Our analysis relies on showing that designing optimization algorithms for time-varying problems can be reformulated as an output regulation problem. By leveraging tools from center manifold theory, we derive necessary and sufficient conditions for exact asymptotic tracking. These results, in turn, pave the way for the development of new optimization algorithms. We demonstrate the effectiveness of the approach through numerical experiments on both synthetic problems and the dynamic traffic assignment problem from traffic control.
@@ 1 Introduction
Time-varying optimization problems play a central role in several scientific domains, as they underpin many important contemporary engineering problems. Examples include training in Machine Learning [1, 2], dynamic signal estimation in Signal Processing [3], trajectory tracking in Robotics [4], system optimization in Industrial Control [5], and much more. Historically, discrete-time algorithms for time-varying optimization have been proposed and studied first since they emerged as a natural extension of their time-invariant counterparts, allowing for cost functions and constraints that may change over time [6]. These approaches build on the classical perspective on optimization, which seeks to construct methods to determine optimizers and consist of iterative procedures implemented on digital devices. In recent years, however, numerous optimization algorithms have been formulated and studied as continuous-time processes [7, 8, 9], mainly due to the wide availability of tools for Ordinary Differential Equations (ODEs) that can facilitate their analysis [9].
@@
Figure 1: Architecture of the gradient-feedback design scheme studied in this work. An optimization algorithm is to be designed (bottom block), having access to gradient evaluations of the loss to be minimized (top right block), and generating a sequence of exploration points $x(t)$ at which the gradient shall be evaluated. The loss function varies with time, where the temporal variability $\theta(t)$ is assumed to be unmeasurable and generated by an exosystem (top left block). Shaded blocks emphasize the presence of dynamics.
Motivated by these recent developments, in this paper, we study time-varying convex optimization problems and focus on the use of continuous-time dynamics to track exactly optimal solutions. Several approaches have already been developed for this purpose [10, 11, 12, 13], yet all these techniques implicitly require full knowledge of the temporal variability of the problem [13]. Unfortunately, in most practical applications, having such knowledge is impractical, either because the temporal variability enters the optimization in the form of exogenous disturbances that are unknown and cannot be measured (see, e.g., [14, 15]), or simply because it is unrealistic to ask for a noiseless model of the problem’s variability. Departing from this, in this paper we pose the following question: is it possible to track (exactly and asymptotically) a minimizer of a time-varying optimization problem without any knowledge of the temporal variability of the optimization? Interestingly, we prove a fundamental result showing that tracking can be achieved if and only if the temporal variability of the problem can be ‘observed’ by the algorithm, and the latter incorporates a suitably reduplicated model of such a variability. We refer to this conclusion as the internal model principle of time-varying optimization, akin to its control-theoretic counterpart [16, 17]. Our approach relies on reinterpreting the optimization algorithm design as a nonlinear, multivariable regulation problem [18], and our analysis uses tools from center manifold theory [19, 17]. Fig. 1 illustrates the architecture of the gradient-feedback design scheme studied in this paper.
@@
Related works. The literature on methods for time-varying optimization is mainly divided into two classes of solutions. The first class consists of methods that do not utilize any model of the temporal variability of the problem [20, 21, 22, 23]; instead, they seek to solve a sequence of static problems. Several established approaches belong to this class, including the online gradient descent method [24] and the online Newton step algorithm [25]—see [26] and references therein. Because the temporal variability of the problem is unknown (or ignored), algorithms in this class can make a decision only after each variation has been observed, thus incurring a certain regret. Mathematically, these techniques are capable of reaching only a neighborhood of an optimizer [13], and exact tracking is out of reach in general for these approaches. This may not be surprising, as it has been recognized recently that these algorithms can be interpreted as basic integral controllers [27]. Moreover, suitable assumptions need to be made to provide convergence guarantees for these methods; we refer to [13, 26] and references therein for a detailed discussion. In contrast, the second class of methods uses a model of the temporal evolution of the problem to exactly track the optimal trajectories [28, 2, 29, 30]. Particularly celebrated is the prediction-correction algorithm (see [10], the recent work [11], and the discrete implementations [31]), whereby, at each time, a prediction step is used to anticipate how the optimizer will evolve over time, and a correction step is used to seek a solution to each instantaneous optimization problem. We refer to [13] for a recent overview of the topic. Recent years have witnessed a growing interest in this problem: [4] uses contraction to study these methods; [32] uses sampling to estimate the temporal variability of the problem; a recent survey has appeared in [33]; constrained optimization problems are studied in [34]; recent works [35, 36, 37, 38] have modified these methods to achieve fixed-time convergence. With respect to this body of literature, which assumes precise knowledge of the temporal variability of the problem (see detailed discussion in Section 2.1), our focus here is to answer the following fundamental question: what is the least-restrictive information that is needed to track, asymptotically with zero error, the optimizer of a time-varying problem? To the best of the authors’ knowledge, a rigorous answer to this question is missing in the literature.
@@
In this work, we recast the problem of designing optimization algorithms for time-varying optimization as a nonlinear output regulation problem. This connects our work with the literature on output regulation, which is a well-established area of research. Initial works can be traced back to the 1970s [16, 39, 18] focusing on linear systems, and later extended to nonlinear systems using both local [40, 41] and global [42] approaches for the analysis. In recent years, the field has received new attention using modern methods; see, e.g., [43, 44, 45] and the recent tutorial [46]. To the best of our knowledge, this is the first work in the literature that establishes a connection between the problem of time-varying optimization and that of output regulation.
@@
Organization. Section 2 presents the problem of interest, Section 3 the parameter-feedback problem (where the temporal variability of the problem can be explicitly measured), and Section 4 the gradient-feedback problem (where the algorithm has access only to first-order function evaluations). Section 5 discusses the tracking accuracy in relationship to the fidelity of the internal model, Section 6 presents extensions to constrained optimization and discrete-time algorithms, Section 7 validates numerically the results, and Section 8 illustrates our conclusions. Finally, in the Appendix, we summarize basic facts on center manifold theory used in the paper.
Notation. We denote by $\mathbb{S}^{n}$ the space of $n\times n$ symmetric real matrices. Given an open set $U,$ we say that $f:U\rightarrow\real$ is of differentiability class $C^{k}$ if it has a $k\textsuperscript{th}$ derivative that is continuous in $U.$ The gradient of $f(x,\theta):\real^{n}\times\Theta\rightarrow\real,\Theta\subseteq\real^{p},$ with respect to ${x}\in\real^{n}$ is denoted by $\nabla_{{x}}f({x},\theta):\real^{n}\times\Theta\rightarrow\real^{n}$. The partial derivatives of $\nabla_{{x}}f({x},\theta)$ with respect to x{x} and $\theta$ are denoted by $\nabla_{{xx}}f({x},\theta):\real^{n}\times\Theta\rightarrow\mathbb{S}^{n}$ and $\nabla_{{x}\theta}f({x},\theta):\real^{n}\times\Theta\rightarrow\real^{n\times p}$, respectively.
@@ 2.1 Optimization objectives
$\min_{x\in\real^{n}}f(x,\theta(t)),$ (1)
where $t\in\mathbb{R}_{\geq 0}$ denotes time and $f:\real^{n}\times\Theta\rightarrow\real,$ $\Theta\subseteq\real^{p},$ is a loss function that is parametrized by the time-varying parameter vector $\theta:\mathbb{R}_{\geq 0}\rightarrow\Theta.$ We make the following standard assumptions on the loss.
@@
The map $x\mapsto f(x,\theta)$ is convex and $x\mapsto\nabla_{x}f(x,\theta)$ is Lipschitz continuous in n, for each $\theta\in\Theta.$ $\square$
@@
There exists a smooth (i.e., $C^{\infty}$) vector field $s:\Theta\rightarrow\real^{p}$ and $\theta(0)\in\Theta$ such that the parameter vector $\theta(t)$ satisfies
@@
For simplicity of the presentation, we assume that $\Theta$ is some neighborhood of the origin of p, corresponding to the requirement that $\theta=0$ is an equilibrium of (2). We put no restrictions on the size of this neighborhood (which is, e.g., allowed to be the entire space $\Theta=\real^{p}$), and thus on the size of $\theta(t)$ nor on its temporal variation. Moreover, there is no restriction with asking that $\theta=0$ is an equilibrium because, if the equilibrium is located at any other point, the latter can be shifted to the origin through a time-invariant change of variables that does not alter the critical points of (1) or the validity of Assumption 1.
@@
In what follows, we say that $x^{\star}(t),$ with $x^{\star}:\mathbb{R}_{\geq 0}\rightarrow\real^{n}$, is a critical trajectory of (1) if it satisfies:
@@
In our analysis, we will use a critical point with $\theta$ at rest, which is a constant vector $x^{\star}_{\circ}\in\real^{n}$, defined implicitly as:
@@
where $\eta>0$ models the algorithm’s step size. It is known [49] that this class of algorithms is capable of computing (asymptotically) critical points of (1) only when the latter are constant; in all other cases, the algorithm converges to a neighborhood of a critical point and convergence is inexact. We refer the reader to Remark 11 for an insightful interpretation of this fundamental limitation. $\square$
@@
where $\nabla_{xx}f(x,\theta)$ is the Hessian of $f(x,\theta)$, and $\nabla_{x\theta}f(x,\theta)$ is the partial derivative of $\nabla_{x}f(x,\theta)$ with respect to $\theta$. Under strong convexity assumptions, this algorithm is known to converge exponentially, with zero asymptotic error, to a critical trajectory [10, 11]. Variations of this algorithm have also been proposed to achieve fixed-time or finite-time convergence [35]. We conclude by noting that implementing (8) requires (i) the cost to be twice continuously differentiable and strongly convex, (ii) knowledge of the maps $s(\theta)$, $\nabla_{x}f(x,\theta)$, $\nabla_{xx}f(x,\theta)$, and $\nabla_{x\theta}f(x,\theta)$, and (iii) knowledge of $\theta(t).$ $\square$
@@ 2.2 Problem statement
In line with the literature [26, 10, 11, 35, 12], we focus on gradient-type algorithms for solving (1), which are algorithms that have access to first-order oracles of the cost; that is, function evaluations of the map:
@@
In the optimization literature [26], function evaluations of (9) are generally obtained through two main approaches: (i) the algorithm has access to both the analytical form of $\nabla_{x}f(x,\theta)$ and $\theta(t)$ at each time $t$ (either because this signal is known or measurable), or (ii) an oracle returns evaluations of (9) (computed, e.g., through numerical differentiation). These alternatives will be discussed more in detail in (O1)–(O4), shortly below. $\square$
We let the optimization algorithm be described by a dynamic internal state $z(t),$ taking values on an open subset $\mathcal{Z}\subseteq\real^{n_{c}},n_{c}\in\mathbb{N}_{>0}$. The optimization algorithm generates a sequence of points $x(t)\in\real^{n}$ (called exploration signal), and has access to function evaluations of (9) at these points (called gradient feedback signal):
@@
where $F_{c}:\mathcal{Z}\times\real^{n}\rightarrow\real^{n_{c}}$ and $G_{c}:\mathcal{Z}\rightarrow\real^{n}$ are functions to be designed. In the remainder, we refer to (10) as a dynamic gradient-feedback optimization algorithm. The architecture of the presented gradient-feedback algorithm is illustrated in Fig. 1. For simplicity of the presentation, we will require that $F_{c}(z,y)$ and $G_{c}(z)$ are such that [^3] Notice that this is without loss of generality since $F_{c}(z,y)$ and $G_{c}(z)$ are to be designed and $x_{\circ}^{\star}$ is known through (6).:
$0$ $=F_{c}(z_{\circ}^{\star},0)\qquad\text{and}\qquad x_{\circ}^{\star}=G_{c}(z_{\circ}^{\star}),$ (11)
for some locally unique $z_{\circ}^{\star}\in\real^{n_{c}}$. Together with (6), this ensures that $z_{\circ}^{\star}$ is an equilibrium of (10) when the gradient feedback signal $y(t)$ is equal to zero.
@@
Table 1: Summary of required knowledge and known guarantees by various established approaches. The symbol $(\checkmark)$ is used when (O1) follows from (E1) and (O2). See Remarks 4, 5, and 7 for discussions. Notice that particular assumptions on the optimization (1) and the exosystem (2) may differ across works.
@@ 3 The parameter-feedback problem
where $H_{c}:\Theta\rightarrow\real^{n}$ is a mapping to be designed. Because of the explicit dependence on $\theta(t),$ we will refer to (14) to as a parameter-feedback optimization algorithm. In analogy with (11), for simplicity of the presentation, in this section, we will impose that $H_{c}(\theta)$ satisfies:
@@
For the algorithm (14) considered in this section, Problems 2.2 and 1 are reformulated, respectively, as follows.